OpenStudy (anonymous):

Please help!!! Ricky is 24 years old and starting an IRA(individual retirement account). He is going to invest $200 at the beginning of each month. The account is expected to earn 2.95% interest, compounded monthly. How much money will Ricky have in his IRA when he retires, at age $65?

OpenStudy (tkhunny):

Do you have an accumulation formula for payments at the beginning of each period? We can build it if you don't have it.

OpenStudy (anonymous):

I don't

OpenStudy (ivettef365):

do you have choices

OpenStudy (anonymous):

Can you help me solve this

OpenStudy (anonymous):

Yes they are... $108,024.51 $125,609.25 $191,398.49 $265,037.90

OpenStudy (ivettef365):

I found this website that is more accurate to your question, because I didn't realize about the monthly deposit of $200

OpenStudy (ivettef365):

It gives me 192,067

OpenStudy (ivettef365):

so the closer to that number is c

OpenStudy (anonymous):

I need help solving it I can't pull the calculator up on my phone

OpenStudy (tkhunny):

You'll need a calculator with an exponent button.

OpenStudy (ivettef365):

now it gives me the 191,398

OpenStudy (ivettef365):

I was adding 200 extra on the investment

OpenStudy (anonymous):

Thanks so much can you help with a few more

OpenStudy (ivettef365):

if you go to that website you get the answer, try it and let me know, on starting amount just put 0

OpenStudy (ivettef365):

yw :)

OpenStudy (tkhunny):

Ricky is 24 years old and starting an IRA(individual retirement account). He is going to invest $200 at the beginning of each month. The account is expected to earn 2.95% interest, compounded monthly. How much money will Ricky have in his IRA when he retires, at age 65? Okay, here we go. 65 - 24 = 41 41 years or 41*12 ==> 492 months. i = 2.95% -- Annual Interest Rate j = I/12 = 0.002458333 -- Monthly Interest Rate r = 1+j -- Monthly Accumulation Factor. We can learn how to do ALL such problems if you can admit to getting that far and then clearing two more small hurdles.